### v3

# A semi-parametric approach for estimating consumer valuation distributions using second price auctions

Abstract
We focus on online second price auctions, where bids are made sequentially, and the winning bidder pays the maximum of the second-highest bid and a seller specified starting price. For many such auctions, the seller does not see all the bids or the total number of bidders accessing the auction, and only observes the current selling prices throughout the course of the auction. We develop a novel semi-parametric approach to estimate the underlying consumer valuation distribution based on this data. Previous semi-parametric or non-parametric approaches in the literature only use the final selling price and assume knowledge of the total number of bidders. The resulting estimate, in particular, can be used by the seller to compute the optimal profit-maximizing price for the product. Our approach is free of tuning parameters, and we demonstrate its computational and statistical efficiency in a variety of simulation settings, and also on an Xbox 7-day auction dataset on eBay.
Second price auction, Semi-parametric maximum likelihood estimation, consumer valuation distribution, standing price sequence,
keywords:
[…]
, , , and
coauco-first authors
1 Introduction
In a second price auction with starting price, the product on sale is awarded to the highest bidder if the corresponding bid is higher than a seller-specified starting price $r$. The price paid by the winner is however, the maximum of the starting price and the second highest bid. These auctions have been the industry standard for a long time, and are attractive to sellers as they induce the bidders to bid their true “private value” for the product, i.e., the maximum price they wish to pay for it. While some platforms have recently moved to first price auctions, the second price auction is still widely used on E-commerce platforms such as eBay, Rokt, and online ad exchanges such as Xandr. The analysis of data obtained from these auctions presents unique challenges. For a clear understanding of these challenges, we first discuss in detail the auction framework, the observed data and the quantity of interest that we want to estimate/extract.
Auction framework: We consider an auction setting where a single product is on sale for a fixed time window $0,\tau]$. The seller sets the starting price $r$, which is used as the current selling price at time 0. Any bidder who arrives subsequently is eligible to place a bid only if his/her bid value is higher than the current selling price at that time. If the bid is placed, the current selling price is updated to the second-highest bid value among the set of all placed bids up to that time, including this latest bid (the starting price is also treated as a placed bid).  [^1] Typically, a small increment (e.g. $\mathdollar 0.01$) is also added to the second highest bid, but this insignificant increment is unlikely to influence bidder’s behaviour, and we ignore it in our analysis for ease of exposition. For example, suppose the current selling price at a given time is $\mathdollar 4$ and the highest placed bid value up to that time is $\mathdollar 5$. If a bidder comes and bids $\mathdollar 3$, the bid will not be placed. If the bidder were to bid $\mathdollar 4.5$, the bid would be placed and the current selling price would be updated to $\mathdollar 4.5$ (since now the second largest placed bid is $\mathdollar 4.5$). If the bidder were to bid $\mathdollar 6$, the bid would be placed and the current selling price would be updated to $\mathdollar 5$. At the end of the auction period, if no bid above the starting price is placed, the item goes unsold, otherwise it is sold to the highest bidder at the selling price at time $\tau$. This final selling price is the second highest placed bid (including the starting price) throughout the course of the auction.
Figure 1: An illustration of a single second price auction. True bid values are generated from a Pareto distribution and starting price is at $\mathdollar 2$. Blue vertical lines are the bid values, black horizontal lines are the current selling prices, and black dots are the time points when the bids are made.
An illustration of a single second price auction: Figure 1 provides a concrete illustration of how a second price auction works. The data for this auction has been simulated from a setting where the bids follow a Pareto distribution with location parameter $3$ and dispersion parameter $100$. The waiting times between bids are generated from an exponential distribution with rate parameter $1$. In total, $100$ bids are generated in a time period $\tau$ of around $115$ minutes. The starting price for the auction is $\mathdollar 2$. In Figure 1, the bid values and the current selling prices during the course of the auction are represented by the blue vertical lines and the black horizontal lines, respectively. The black dots on the black horizontal lines represent the time points (in minutes) of the $100$ bids. As can be seen from Figure 1, the initial selling price is equal to the starting price ($\mathdollar 2$). We have the first bid of around $\mathdollar 8.05$ at $0.55$ minutes. Since it’s higher than the starting price of $\mathdollar 2$, this bid will be placed, but the starting price still remains the current selling price at $0.55$ minutes. Recall that arriving consumers only see the current selling price, and hence a consumer with a $\mathdollar 5.09$ valuation would be both willing and eligible to place their bid when the current selling price is $\mathdollar 2$. Hence, the second bid of $\mathdollar 5.09$ occurring at $5.96$ minutes ($5.41$ minutes after the first bid’s occurrence) is placed, and subsequently the current selling price jumps to $\mathdollar 5.09$ as it’s the second highest value among the starting price ($\mathdollar 2$) and the two existing placed bids ($\mathdollar 8.05$ and $\mathdollar 5.09$). We don’t observe any jumps in the current selling price for the next few bids as they all are less than the current selling price of $\mathdollar 5.09$ (and hence are not eligible to be placed). Then we see another jump at $9.65$ minutes, where a bid of $\mathdollar 12.82$ which makes the current selling price jump to $\mathdollar 8.05$. The next few bids again happen to be less than the current selling price ($\mathdollar 8.05$). At around $24$ minutes, we see a last jump in the current selling price to $\mathdollar 10.14$ (based on a new bid of $\mathdollar 10.14$ which exceeds $\mathdollar 8.05$). The subsequent bids are all less than $\mathdollar 10.14$, it remains the current selling price throughout the rest of the auction period. This can be observed through the flat horizontal black line at $\mathdollar 10.14$ in the time period of $(24.7,114.43)$ minutes. The final selling price for the auction is therefore $\mathdollar 10.14$. The observed data for the above auction is the sequence of current selling prices given by $(\mathdollar 2,\mathdollar 5.09,\mathdollar 8.05,\mathdollar 10.14)$ and the sequence of times at which there was a change in the current selling price, given by $(5.96,9.65,24)$.
Observed data: The observed data is the sequence of current selling price values (also sometimes referred to as the standing price) throughout the course of the auction, and the times at which there is a change in the selling/standing price. Typically, such data is available for multiple auctions of the same product. For example, in Section 5, we analyze data with current selling prices for $93$ different eBay $7$-day auctions for Xbox. A key observation to make here is that consumers who access the auction but have bids which are less than the current selling price (standing price) are not allowed to place their bids. In other words, instead of observing the bids of all the customers who access the auction, we only observe the running second largest value of such bids.
Figure 2: An illustration of 4 separate second price auctions for a given product. For each auction, true bid values are generated from a discrete Uniform distribution and starting price is respectively set to $\mathdollar 10,\mathdollar 5,\mathdollar 13,\mathdollar 17$. Blue vertical lines are the bid values, black horizontal lines are the current selling prices, and black dots are the time points when the bids are made.
An illustration of multiple second price auctions: Figure 2 provides a concrete illustration of a setting with multiple (four) second price auctions for the same product. The data for these auctions has been simulated from a setting where the bids follow a discrete uniform distribution with lower bound 1 and upper bound 25. The waiting times between bids are generated from an exponential distribution with rate parameter $1$, and the time period for each auction is 10 days. The starting prices for the four auctions, respectively, are $\mathdollar 10,\mathdollar 5,\mathdollar 13,\mathdollar 17$. Just like in Figure 1, the bid values and the current selling prices during the course of each auction are represented by the blue vertical lines and the black horizontal lines, respectively. The black dots on the black horizontal lines represent the time points (in days) of the bids. We can see that for auctions in the upper left and upper right, the corresponding item is sold above the starting price (at final selling price of $\mathdollar 19$ and $\mathdollar 25$ respectively). On the other hand, for the auction in the lower left, the item is sold at the starting price, since there is only one bid placed above/at the starting price. Finally, for the auction in the lower right, the item is not sold since no bid is placed above or at the starting price. We will keep coming back to this integrated example to clarify and illustrate various concepts and notions that are introduced in the paper.
Quantity of interest: Each bidder in the consumer population is assumed to have an independent private valuation (IPV) of the product. The IPV assumption in particular makes sense for products that are used for personal use/consumption (such as watches, jewelry, gaming equipment etc.) and is commonly used in the modeling of internet auctions (see Song (2004); Hou and Rego (2007); George and Hui (2012) and the references therein). Economic theory suggests that the dominant strategy for a bidder in a second price auction is to bid one’s true valuation (Vickrey (1961)). The quantity that we want to estimate from the above data is the distribution of the valuations of the product under consideration for the consumer population. We refer to this as the consumer valuation distribution, and denote it by $F$. As noted in George and Hui (2012), knowledge of $F$ provides the consumer demand curve for the product, and hence can be used by the seller to identify the profit-maximizing price (see the discussion in Section 4.1 from George and Hui (2012)).
The problem of demand-curve/valuation distribution estimation using auction data has been tackled in the last two decades for a variety of auction frameworks, see Song (2004); Park and Bradlow (2005); Bradlow and Park (2007); Yao and Mela (2008); George and Hui (2012); Backus and Lewis (2025) and the references therein. Some of these papers assume a parametric form for $F$, but as George and Hui (2012) argue, available auction data may often not be rich enough to verify the validity of the underlying parametric forms. George and Hui (2012) consider the second-price ascending bid auction framework for a single (homogeneous) product described above, and develop a Bayesian non-parametric approach to estimate $F$ using only the final selling prices in multiple auctions for a jewelry item. Note that the final selling price is not only the second highest placed bid in the auction period, it is also the second maximum order statistic of the (potential or placed) bids of all consumers who access the auction. Using only the final selling prices leads to identifiability problems, and George and Hui (2012) address this by assuming that the total number of consumers accessing the auction is also known. This information is available for the particular jewelry dataset from a third-party vendor, but is generally not available for most such datasets. When the total number of consumers accessing the auction is not known, the identifiability issue can be solved by using another order statistic (Song (2004)) such as the largest placed bid value (if available) along with the final selling price. (Backus and Lewis, 2025, Section 5) use such an approach for analyzing a dataset containing compact camera auctions on eBay.
Paper Auc. type Statistical Paradigm Utilizes entire standing price sequence Needs information on unobserved bidders or bid values Allows multiple bidding and sniping in same auction Allows buy-it-now option
Bradlow & Park (2007) $1^{st}$ Parametric (Bayesian) Yes No Yes No
Chan et al. (2007) $1^{st}$ Parametric Yes No Yes Yes
Song (2004) $2^{nd}$ Semi and nonparametric No (only uses $2^{nd}$ and $3^{rd}$ highest price) No Yes No
Yao & Mela (2008) $2^{nd}$ Parametric (Bayesian) Yes Yes (needs highest bid value in the auction) Yes No
George & Hui (2012) $2^{nd}$ Non-parametric (Bayesian) No (only uses final selling price) Yes (needs total number of participating bidders) Yes No
Backus & Lewis (2024) $2^{nd}$ Non-parametric No (only uses $1^{st}$ and $2^{nd}$ highest price) Yes (needs highest bid value in the auction) Yes No
This paper $2^{nd}$ Semi-parametric Yes No No No
Table 1: A tabular comparison of the main features of some key related papers and the proposed methodology.
Table 1 provides a comparison of the main relevant features of key related papers and the proposed methodology in the paper. As the table demonstrates (and to the best of our knowledge), none of the existing non-parametric/semi-parametric methods for second price auctions use the entire sequence of current selling price values to estimate the consumer valuation distribution $F$. This was the primary motivation for our work, as only using the final selling price (and possibly the maximum placed bid, if available) leaves out a lot of available information. Including the current selling price information throughout the course of the auction however, involves significant conceptual and methodological challenges. As observed previously, the final selling price (second largest placed bid) is also the second largest (potential or placed) bid of all the consumers who accessed the auction. Hence, under relevant assumptions (see beginning of Section 2), the final selling price can be interpreted as the second largest order statistic of i.i.d. samples from $F$. This interpretation is central to the methodology developed in George and Hui (2012) and Song (2004). However, as the authors in George and Hui (2012) point out, the second largest current selling price throughout the course of the auction is not necessarily the third largest order statistic among all (placed or potential) bids, unless some severely restrictive and unrealistic assumptions are imposed on the order in which bidders arrive in the auction. Hence, extending the methodology in George and Hui (2012) to include the entire sequence of current selling prices is not feasible; a completely new method of attack is needed.
In this paper, we fill this gap in the literature and develop novel methodology for semi-parametric estimation of the consumer valuation distribution in second-price ascending bid auctions which uses the entire sequence of current selling prices. Additionally, the total number of consumers accessing the auction is not assumed to be known (unlike George and Hui (2012)), and the highest placed bid in the auction is also not assumed to be known (unlike Backus and Lewis (2025)). Incorporating the above novel features is very challenging, and the methodological development (provided in Section 2) is quite involved. The extensive simulation results in Section 4 demonstrate the significant accuracy gains in the estimation of $F$ that can be obtained by using the entire sequence of current selling prices as opposed to just the final selling price. We note that incorporating the current selling prices during the entire auction does come with a cost. In particular, two additional assumptions (compared to those in George and Hui (2012)) that need to be made about the rate of bidder arrival and number of bids made by a single consumer. We find some deviations from these two assumptions in the eBay $7$-day Xbox data that we analyze in Section 5, but these deviations are minor (see discussion at beginning of Section 2). In such settings, it is reasonable to expect that the advantage of incorporating substantial additional data/information outweighs the cost of these approximations/deviations. We conclude the paper with a discussion of future directions in Section 6.
2 Methodology for learning the consumer valuation distribution $F$
We start by discussing the main assumptions needed for the subsequent methodological development. We consider a setting where we have data from several independent, non-overlapping auctions for a single (homogeneous) product. As mentioned previously, we work within the IPV framework which is quite reasonable for items/products that are used for personal consumption. Similar to George and Hui (2012), we will assume that the collection of bidders who access the auction is an i.i.d. sample from the consumer population for the corresponding product, and that the collection of private product valuations of these bidders is an i.i.d. sample from the consumer valuation distribution $F$. As stated in the introduction, it can be shown that the dominant strategy for a bidder in a second price auction is to bid his/her true valuation. Under this strategy, any consumer who accesses the auction would bid his/her valuation with no need for multiple bidding, and we assume this behavior. We do note that in practice, some consumers do not follow this strategy. For example, eBay (1995) provides an option called proxy bidding or automatic bidding which allows the computer to automatically place multiple incremental bids below a cutoff price on behalf of the consumer (see also Ockenfels and Roth (2006); P. Bajari and A. Hortacsu (2003)). Since George and Hui (2012) only use the final selling price, they use a weaker assumption which allows for multiple bidding and stipulates that a consumer bids his/her true valuation sometime before the end of the auction.
Our final assumption is regarding the arrival mechanism of bidders in the auction. We assume that consumers arrive at/access the auction according to a Poisson process with constant rate $\lambda$. Again, a rational bidder (based on economic theory) in a second price auction should be indifferent to the timing of his/her bid (see Milgrom (2004); Barbaro and Bracht (2021)), and this assumption makes sense in such settings. Again, we note that late-bidding (sniping) has been observed in some eBay auctions (see Bose and Daripa (2017); Barbaro and Bracht (2021)).
To summarize, our assumptions are identical to those in George and Hui (2012) with the exception of the single bidding-assumption and the constant rate of arrival assumption. While these assumptions are supported in general by economic theory, deviations from these two assumptions have been observed in some online auctions. However, for datasets such as the Xbox dataset analyzed in Section 5, these deviations are minor/anomalies. For example, in the Xbox dataset, only around $10\%$ of the bidders with placed bids show a multiple bidding behavior. In such settings, there is certainly value in using the subsequent methodology which uses the entire current selling price/standing price profile and does not assume knowledge of the total number of consumers accessing the auctions. If there is strong evidence/suspicion that the assumptions are being extensively violated, of course the results from this methodology should be treated with due skepticism and caution.
The subsequent methodological development in this section is quite involved, and we have tried to make it accessible to the reader by dividing it into subsections based on the major steps, and then highlighting the key milestones within each subsection, wherever necessary. We start by finding the joint density of the observed data obtained from a single second price auction.
2.1 Joint density of the observed data in a single second price auction
Consider a given (single) second price auction with starting price $r$. Hence, the initial selling price, denoted by $X_{0}$, is equal to $r$. The first time a consumer with bid value greater than $r$ arrives at the auction, that bid is placed but the current selling price remains $r$. Subsequently, the current selling price (standing price) changes whenever a bid greater than the existing selling price is placed. Let $M$ denote the number of times the selling price changes throughout the course of the auction. When $M>0$, let $\{X_{i}\}_{i=1}^{M}$ denote the sequence of current selling prices observed throughout the course of the auction, with $X_{i}$ denoting the new selling/standing price after the $i^{th}$ change. When $M>0$, let $T_{i}$ denote the intermediate time between the $i^{th}$ and $(i+1)^{th}$ changes in the selling/standing price for $0\leq i\leq M-1$. In particular, it follows that when $M>0$, $T_{0}$ denotes the waiting time from the start of the auction until the moment when for the first time, the selling price changes to a higher value than the starting price $r$. When $M=0$, we define $T_{0}=\tau$. Finally, let $O$ be a binary random variable indicating whether the item is sold before the end of the auction, i.e., $O=1$ indicates the item is sold and $O=0$ indicates that the item is not sold. Our observed data comprises of $M$, $O$, $\{X_{i}\}_{i=1}^{M}$, and $\{T_{i}\}_{i=0}^{M-1}$.
We define $T_{M}=\tau-\sum_{i=0}^{M-1}T_{i}$ as the time after the last selling price change and until the auction closes. As discussed earlier, the number of consumers/bidders accessing a given (single) second price auction, denoted by $N$, remains unobserved in our setup. Based on our assumption regarding the arrival mechanism (Poisson process with constant rate of arrival $\lambda$) of bidders in the auction, we note that $N\sim\text{Poisson}(\lambda\tau)$.
Note that there are three scenarios at the end of the auction: (a) the item is sold above the starting price ($M>0,O=1$), (b) the item is sold at the starting price ($M=0,O=1$), and (c) the item is not sold ($M=0,O=0$). The following lemma provides a unified formula for the joint density of the observed data encompassing all these three scenarios.
Lemma 2.1.
For the second price auction described above, the joint density of $M$, $O$, $\{X_{i}\}_{i=1}^{M}$, $\{T_{i}\}_{i=0}^{M-1}$ at values $m$, $o$, $\{x_{i}\}_{i=1}^{m}$, $\{t_{i}\}_{i=0}^{m-1}$ is given by
$\begin{split}{}&g\Big(m,o,\{x_{i}\}_{i=1}^{m},\{t_{i}\}_{i=0}^{m-1}\Big)\\ ={}&\exp(-\lambda\tau)\ 2^{m}\ \Big(\lambda^{m+1}\ t_{0}\ \big(1-F(x_{m})\big)\Big)^{1_{\{o=1\}}}\exp\bigg(\lambda\sum_{i=0}^{m}F(x_{i})t_{i}\bigg)\bigg(C_{1}\prod_{i=1}^{m}f(x_{i})\bigg)^{1_{\{m>0\}}},\end{split}$
where $C_{1}$ does not depend on $F$, $x_{0}$ represents the starting price ($r$), $\lambda$ denotes the constant rate of arrival of the bidders throughout the course of the auction, and $f$ represents the density function corresponding to $F$.
The proof of Lemma 2.1 is quite involved and is provided in the supplementary document. Next, we generalize our analysis to consider data from several independent, non-overlapping auctions of identical copies of a single (homogeneous) product.
2.2 Likelihood based on the observed data from multiple identical, non-overlapping second price auctions
Suppose we consider $K$ independent second price auctions of identical copies of an item (with possibly different starting prices $r_{1},r_{2},\ldots,r_{K}$). The observed data is $\{(M_{k},O_{k})\ ,\ \{(X_{i,k}\ ,\ T_{i-1,k})\}_{i=1}^{M_{k}}\}_{k=1}^{K}$, where $M_{k}$ denotes the number of selling price changes for the $k^{th}$ auction, $O_{k}$ denotes the indicator of the item being sold at the end of the $k^{th}$ auction, $\{X_{i,k}\}_{i=1}^{M_{k}}$ denote the selling/standing price sequence for the $k^{th}$ auction, and $\{T_{i-1,k}\}_{i=1}^{M_{k}}$ denote the sequence of intermediate waiting times between successive changes in the standing prices. Finally, let $T_{M_{k},k}=\tau-\sum_{i=0}^{M_{k}-1}T_{i,k}$, for all $k=1,2,\ldots,K$. Since the auctions are independent, it follows by Lemma 2.1 that the likelihood function of the unknown parameters $\lambda$ and $F$ for observed data values $\{(m_{k},o_{k})\ ,\ \{(x_{i,k}\ ,\ t_{i-1,k})\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ is given by $Lik(\lambda,F)=\prod_{k=1}^{K}g(m_{k},o_{k},\{x_{i,k}\}_{i=1}^{m_{k}},\{t_{i,k}\}_{i=0}^{m_{k}-1})$.
Ideally, one would like to obtain estimates of $\lambda$ and $F$ by maximizing the function $Lik(\lambda,F)$. However, this likelihood function is intractable for direct maximization. A natural direction to proceed is to use the alternative maximization approach, which produces a sequence of iterates by alternatively maximizing $Lik$ with respect to $F$ given the current value of $\lambda$ and then maximizing $Lik$ with respect to $\lambda$ given the current value of $F$. Especially the maximization with respect to $F$ (given $\lambda$) requires intricate analysis and careful reparametrization, and we describe the details in Sections 2.3 and 2.4 below.
2.3 Estimation of $F$ given $\lambda$: Some new notation based on pooled standing prices across all auctions
Note that the function $F$ is constrained to be non-decreasing. A key transformation to a constraint-free parametrization (described in Section 2.4 below) is needed to facilitate the conditional maximization of $Lik$ with respect to $F$. A crucial precursor to this re-parametrization is introduction of some new notation obtained by merging the standing prices from all the $K$ different auctions together. Let $L=\sum_{k=1}^{K}M_{k}$ denote the total number of standing/selling price changes in all the $K$ auctions. Recall that $\{(m_{k},o_{k})\ ,\ \{(x_{i,k}\ ,\ t_{i-1,k})\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ denote the observed data values, and $t_{m_{k},k}=\tau-\sum_{i=0}^{m_{k}-1}t_{i,k}$ for $1\leq k\leq K$. Let $\ell$ denote the observed value of $L$. We will denote by ${\bf\bar{x}}=(\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\ell})$ the arrangement/ordering of the pooled collection $\{\{x_{i,k}\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ such that $\bar{x}_{1}<\bar{x}_{2}<\ldots<\bar{x}_{\ell}$; under the assumption that $F$ is a continuous cdf, there should be no ties in the entries of $\bf\bar{x}$} with probability one. In other words, we pool the standing prices from all the auctions (excluding the starting prices) and then arrange them in ascending order as $(\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\ell})$. Also, for $1\leq i\leq L$, we define $\bar{t}_{i}=t_{\bar{i},\bar{k}}$ where $\bar{i}$ and $\bar{k}$ are such that $\bar{x}_{i}=x_{\bar{i},\bar{k}}$. Further, let
[…]
item is sold above the starting price,
[…]
Example (continued): Consider the example in Figure 2, where we have a setting with $K=4$ independent second price auctions, with starting prices $(r_{1},r_{2},r_{3},r_{4})=(10,5,13,17)$. The first auction has $m_{1}=3$ standing price changes with $(x_{1,1},\ x_{2,1},\ x_{3,1})=(12,15,19)$, the second auction has $m_{2}=4$ standing price changes with $(x_{1,2},\ x_{2,2},\ x_{3,2},\ x_{4,2})=(16,18,20,25)$, the item is sold at the starting price $r_{3}=13$ in the third auction ($m_{3}=0,o_{3}=1$), and the item is unsold in the fourth auction ($m_{4}=0,o_{4}=0$). Pooling and rearranging the standing prices (excluding starting prices) from all the auctions, we see that $\ell=3+4+0+0=7$, and
${\bf\bar{x}}=(\bar{x}_{1},\bar{x}_{2},\bar{x}_{3},\bar{x}_{4},\bar{x}_{5},\bar{x}_{6},\bar{x}_{7})=(12,15,16,18,19,20,25).$
Note that the auction item is sold above the starting price in the first two auctions, and the final selling prices are $x_{m_{1},1}=19$ and $x_{m_{2},2}=25$ respectively. Examining the positions of these two prices in $\bf\bar{x}$} gives us $S=\{5,7\}$. Finally, $K_{s}=\{1,2,3\}$ is the collection of auction indices where the item is sold.
Using the newly introduced notation above and Lemma 2.1, it follows that the likelihood function is given by
$Lik\left(\lambda,F\right)={}$ $C^{\star}\exp\big(-K\lambda\tau\big)\ \lambda^{\ell+\left|K_{s}\right|}\ 2^{\ell}\prod_{k\in K_{s}}t_{0,k}\ \prod_{i\in S}\big(1-F(\bar{x}_{i})\big)\bigg[\prod_{j=1}^{\ell}f(\bar{x}_{j})\bigg]^{1_{\{\ell>0\}}}$
${\color[rgb]{0,0,0}\times\prod_{k=1}^{K}(1-F(r_{k}))^{1_{\{}m_{k}=0,o_{k}=1\}}\exp\bigg(\lambda\bigg(\sum_{k=1}^{K}F(r_{k})t_{0,k}+\sum_{j=1}^{\ell}F(\bar{x}_{j})\bar{t}_{j}\bigg)\bigg)},$ (2.1)
where $C^{\star}$ doesn’t depend on $(\lambda,F)$. Maximization of the above likelihood (given $\lambda$) over absolutely continuous CDFs leads to one of the standard difficulties in non-parametric estimation. As $F$ moves closer and closer to a CDF with a jump discontinuity at any $\bar{x}_{j}$, the function $Lik(\lambda,F)$ converges to infinity for every fixed $\lambda$. Hence, any absolutely continuous CDF with a density function cannot be a maximizer of the above profile likelihood function. Following widely used convention in the literature (see Murphy (1994), Vardi (1982)), we will extend the parameter space to allow for the MLE of $F$ to be a discrete distribution function. To allow for discrete CDFs, we replace $f(\bar{x}_{j})$ by $F(\bar{x}_{j})-F(\bar{x}_{j}-)$. Thus the adapted likelihood can be written as
$Lik_{A}\left(\lambda,F\right)$ (2.2)
$=$ $C^{\star}\exp\big(-K\lambda\tau\big)\ \lambda^{\ell+\left|K_{s}\right|}\ 2^{\ell}\ \big(\prod_{k\in K_{s}}t_{0,k}\big)\ \prod_{i\in S}\big(1-F(\bar{x}_{i})\big)\prod_{k=1}^{K}\big(1-F(r_{k})\big)^{1_{\{}m_{k}=0,o_{k}=1\}}$
$\times\exp\bigg(\lambda\bigg(\sum_{k=1}^{K}{\color[rgb]{0,0,0}F(r_{k})}t_{0,k}+\sum_{j=1}^{\ell}F(\bar{x}_{j})\bar{t}_{j}\bigg)\bigg)\ \bigg[\prod_{j=1}^{\ell}\Big(F(\bar{x}_{j})-F(\bar{x}_{j}-)\Big)\bigg]^{1_{\{\ell>0\}}},$
where $\bar{x}_{0}=0$. We now establish a final bit of notation necessary to introduce the constraint-free reparametrization of $F$. We now pool the $\ell+K$ standing prices from all the auctions (including the starting prices), i.e., $\{\{x_{i,k}\}_{i=0}^{m_{k}}\}_{k=1}^{K}$, and arrange them in ascending order as $z_{1}<z_{2}<\cdots<z_{\ell+K}$, and denote ${\bf z}:=(z_{1},z_{2},\cdots,z_{\ell+K})$. Under the assumption that $F$ is a continuous cdf, there should be no ties in the entries of $\bf\bar{x}$} with probability one. The only other entries in $\bf z$} are the $K$ starting prices. In practice, it is possible that there are ties in the starting prices, in which case we add a very small noise to the starting prices to ensure that there are no ties in the entries of $\bf z$}. Similar to $\bar{t}_{i}$, we define $\widetilde{t}_{i}=t_{\widetilde{i},\widetilde{k}}$ where $\widetilde{i}$ and $\widetilde{k}$ are such that $z_{i}=x_{\widetilde{i},\widetilde{k}}$. Further, let
[…]
item is sold at or above the starting price,
$\bf{u}:={}$ $\{u_{1},u_{2},\ldots,u_{\ell}\},\ \text{where }u_{i}=\ \text{position of}\ \bar{x}_{i}\ \text{in}\ {\bf z},\ \text{i.e., }\bar{x}_{i}=z_{u_{i}}.$
Since the entries of $\bf\bar{x}$} and $\bf z$} both are arranged in ascending order, it follows that $u_{1}<u_{2}<\ldots<u_{\ell}$.
Example (continued): In the example considered earlier in this subsection with $K=4$ auctions, by pooling the $4$ starting price values with entries of $\bf\bar{x}$} and rearranging them in ascending order, we obtain
${\bf z}=(5,10,12,13,15,16,17,18,19,20,25).$
Recall that the item is sold above the starting price only in the first two auctions. By identifying the positions/ranks of $x_{m_{1},1}$, $x_{m_{2},2}$ and $x_{m_{3},3}$ (with $m_{3}=0$) in $\bf z$}, we obtain $\bar{S}=(4,9,11)$. Similarly, by identifying the positions/ranks of the entries of $\bf\bar{x}$} in $\bf z$}, we obtain ${\bf u}=(3,5,6,8,9,10,11)$.
It is clear from (2.2) that for maximizing $Lik_{A}$ it is enough to search over the class of CDFs with jump discontinuities at elements of $\bf\bar{x}$}, since all other CDFs will have a $Lik_{A}$ value of zero. The next lemma (proof provided in the Supplement) shows that the search for a maximizer can be further restricted to a certain class of CDFs with possible jump discontinuities at elements of $\bf z$}.
Lemma 2.2.
[…]
For any $F\in\mathcal{F}_{\bf z}$, note that $Lik_{A}\left(\lambda,F\right)$ depends on $F$ only through
$\{F(z_{1}),F(z_{2})-F(z_{2}-),\ldots,F(z_{L+K})-F(z_{\ell+K}-)\}$
or equivalently through
$F({\bf z})=(F(z_{1}),F(z_{2}),\cdots,F(z_{\ell+K}))$
(since $F$ only has jump discontinuities at elements of $\bf z$} and is otherwise piece-wise constant). This is typical in a non-parametric setting, and we can hope/expect to only obtain estimates of the valuation distribution $F$ at the observed standing prices (including the starting prices).
2.4 Estimation of $F$ given $\lambda$: A constraint-free reparametrization
Note that the entries of the vector $F({\bf z})$ are non-decreasing, and this constraint complicates the maximization of $F\mapsto Lik_{A}\left(\lambda,F\right)$. So, we transform $F\left(\bf{z}\right)$ to another $\ell+K$ dimensional parameter vector $\boldsymbol{\theta}:=\left(\theta_{1},\theta_{2},\ldots,\theta_{\ell+K}\right)^{T}$ as follows:
$\theta_{i}=\dfrac{G(z_{i})}{G(z_{i-1})}\ ,\ \forall\ 1\leq i\leq(\ell+K),$ (2.3)
where
$G(z_{i})=1-F(z_{i})\ ,\ \forall\ 1\leq i\leq(\ell+K),\ \text{and}\ G(z_{0})=1\ \text{with}\ z_{0}=0.$ (2.4)
Since $F$ is non-decreasing, and takes values in $[0,1]$, it follows that $\theta_{i}\in[0,1]$ (with the convention $0/0:=0$). Focusing our search on the class of CDFs in $\mathcal{F}_{\bf z}$ leads to additional constraints. Since any cdf $F$ in this class has a jump discontinuity at each $\bar{x}_{l}=z_{u_{l}}$, it follows that $G(z_{i})=1-F(z_{i})>0$ for $i<u_{\ell}$, and $G(z_{i})<G(z_{i-1})$ for every $i\in{\bf u}$. In other words, we have $\theta_{i}<1$ for $i\in{\bf u}$, and $\theta_{i}>0$ for $i<u_{\ell}$. There are no other constraints on the elements of $\boldsymbol{\theta}$. Also, we can retrieve $F\left({\bf z}\right)$ given $\boldsymbol{\theta}$ using the following equality.
$F\left(z_{i}\right)=1-G(z_{i})=1-\prod_{j=1}^{i}\theta_{j}\ ,\ \forall\ 1\leq i\leq(\ell+K).$ (2.5)
Now, using (2.2), (2.4) and (2.3), we can rewrite the ‘adapted’ likelihood $Lik_{A}$ in terms of $\boldsymbol{\theta}$ as follows:
$Lik_{A}(\lambda,\boldsymbol{\theta})$
$={}$ $C^{\star}\exp\big(-K\lambda\tau\big)\ \lambda^{\ell+\left|K_{s}\right|}\ 2^{\ell}\prod_{k\in K_{s}}t_{0,k}\prod_{i\in S}G(\bar{x}_{i})\prod_{k=1}^{K}G(r_{k})^{1_{\{}m_{k}=0,o_{k}=1\}}$
$\times\exp\bigg(\lambda\bigg(\sum_{k=1}^{K}\big(1-G(r_{k})\big)t_{0,k}+\sum_{l=1}^{\ell}\big(1-G(\bar{x}_{l})\big)\bar{t}_{l}\bigg)\bigg)\ \bigg[\prod_{l=1}^{\ell}\Big(G(\bar{x}_{l}-)-G(\bar{x}_{l})\Big)\bigg]^{1_{\{\ell>0\}}}$
$={}$ $C^{\star}\exp\big(-K\lambda\tau\big)\ \lambda^{\ell+\left|K_{s}\right|}\ 2^{\ell}\prod_{k\in K_{s}}t_{0,k}\prod_{i\in S}G(\bar{x}_{i})\prod_{k=1}^{K}G(r_{k})^{1_{\{}m_{k}=0,o_{k}=1\}}$
$\times\exp\bigg(\lambda\bigg(K\tau-\sum_{k=1}^{K}G(r_{k})t_{0,k}-\sum_{l=1}^{\ell}G(\bar{x}_{l})\bar{t}_{l}\bigg)\bigg)\ \bigg[\prod_{l=1}^{\ell}\Big(G(\bar{x}_{l}-)-G(\bar{x}_{l})\Big)\bigg]^{1_{\{\ell>0\}}}$
$={}$ $C^{\star}\ \lambda^{\ell+\left|K_{s}\right|}\ 2^{\ell}\prod_{k\in K_{s}}t_{0,k}\prod_{i\in S}G(\bar{x}_{i})\prod_{k=1}^{K}G(r_{k})^{1_{\{}m_{k}=0,o_{k}=1\}}$
$\times\exp\bigg(-\lambda\bigg(\sum_{k=1}^{K}G(r_{k})t_{0,k}+\sum_{l=1}^{\ell}G(\bar{x}_{l})\bar{t}_{l}\bigg)\bigg)\ \bigg[\prod_{l=1}^{\ell}\Big(G(\bar{x}_{l}-)-G(\bar{x}_{l})\Big)\bigg]^{1_{\{\ell>0\}}}$
$={}$ $C^{\star}\ \lambda^{\ell+\left|K_{s}\right|}\ 2^{\ell}\prod_{k\in K_{s}}t_{0,k}\prod_{i\in\bar{S}}G(z_{i})\ \exp\bigg(-\lambda\sum_{i=1}^{\ell+K}G(z_{i})\widetilde{t}_{i}\bigg)\ \bigg[\prod_{l=1}^{\ell}\Big(G(z_{u_{l}}-)-G(z_{u_{l}})\Big)\bigg]^{1_{\{\ell>0\}}},$
where $u_{0}=0$. Using (2.5), we get
$Lik_{A}(\lambda,\boldsymbol{\theta})={}$ $C^{\star}\ \lambda^{\ell+\left|K_{s}\right|}\ 2^{\ell}\bigg(\prod_{k\in K_{s}}t_{0,k}\bigg)\bigg(\prod_{i\in\bar{S}}\prod_{j=1}^{i}\theta_{j}\bigg)\ \exp\bigg(-\lambda\sum_{i=1}^{\ell+K}\widetilde{t}_{i}\bigg(\prod_{j=1}^{i}\theta_{j}\bigg)\bigg)$
$\times\bigg[\prod_{l=1}^{\ell}\bigg((1-\theta_{u_{l}})\prod_{j=1}^{u_{l}-1}\theta_{j}\bigg)\bigg]^{1_{\{\ell>0\}}}.$ (2.6)
A straightforward argument shows that
$Lik_{A}(\lambda,(\theta_{1},\theta_{2},\ldots,\theta_{\ell+K}))=Lik_{A}(c\lambda,(c^{-1}\theta_{1},\theta_{2},\ldots,\theta_{\ell+K}))$
for any constant $c>0$. To address this identifiability issue, we note that no data is observed below $z_{1}$, which is the smallest starting price among all auctions in the data. Hence, it is quite reasonable to assume that $F(z_{1})=0$, which is equivalent to setting $\theta_{1}=1$.
With the identifiability issue resolved, the goal now is to maximize $Lik_{A}$ with respect to $\boldsymbol{\theta}$, where each entry of $\boldsymbol{\theta}$ is in $[0,1]$, $\theta_{i}<1$ for $i\in{\bf u}$, and $\theta_{i}>0$ for $i<u_{\ell}$. We achieve this using the coordinate-wise ascent algorithm. The details of this algorithm are derived in Section 3.
3 Maximizing $Lik_{A}(\lambda,\boldsymbol{\theta})$: Coordinate ascent algorithm
Applying natural logarithm on both sides of the equation in (2.6), we get
$\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))={}$ $\ln(C^{\star})+\big(\ell+\left|K_{s}\right|\big)\ln(\lambda)+\ell\ln(2)+\sum_{k\in K_{s}}\ln(t_{0,k})+\sum_{i\in\bar{S}}\sum_{j=1}^{i}\ln(\theta_{j})$
$-\lambda\sum_{i=1}^{\ell+K}\widetilde{t}_{i}\bigg(\prod_{j=1}^{i}\theta_{j}\bigg)+1_{\{\ell>0\}}\bigg[\sum_{l=1}^{\ell}\ln(1-\theta_{u_{l}})+\sum_{l=1}^{\ell}\sum_{j=1}^{u_{l}-1}\ln\theta_{j}\bigg],$ (3.1)
where $u_{0}=0$. We now introduce notation which allows for a more compact and accessible representation of $\ln(Lik_{A})$. Recall that $\bf z$} is obtained by pooling all the $K$ starting prices and the $\ell=\sum_{k=1}^{K}m_{k}$ ‘non-starting’ standing prices (elements of $\bf\bar{x}$}), and $u_{j}$ represents the position of the $\bar{x}_{i}$ in $\bf z$} for $1\leq j\leq\ell+K$. In particular, $u_{\ell}$ is the position of $\bar{x}_{\ell}$, the largest ‘non-starting’ standing price across all the $K$ auctions in $\bf z$}. In other words, $\bar{x}_{\ell}=z_{u_{\ell}}$. It is possible that $u_{\ell}<\ell+K$. For example, in settings where the starting price in one of the auctions where the item is unsold is larger than $\bar{x}_{\ell}$, it follows that $\bar{x}_{\ell}$ is not the largest entry in $\bf z$} and $u_{\ell}<\ell+K$. With this background, we define
$l_{i}$ $=0\quad\text{if}\quad 1\leq i\leq u_{1}-1,$
$l_{i}$ $=j\quad\text{if}\quad u_{j}\leq i\leq u_{j+1}-1\ ,\ \text{for}\ i=u_{1},u_{1}+1,\ldots,u_{\ell}-1,$
$\text{and}\ l_{i}$ $=\ell\quad\text{if}\quad u_{\ell}\leq i\leq\ell+K.$ (3.2)
In other words, note that $u_{1}<u_{2}<\cdots<u_{\ell}$ induce an ordered partition of the set $\{1,2,\cdots,u_{\ell}-1\}$ into $\ell$ disjoint subsets
$\{1,\cdots,u_{1}-1\},\{u_{1},\cdots,u_{2}-1\},\cdots,\{u_{\ell-1},\cdots,u_{\ell}-1\}.$
Hence, any $1\leq i\leq u_{\ell}$ belongs to one of the subsets in the above partition, and $l_{i}$ is defined to be one less than the position of that subset in the partition. For $u_{\ell}\leq i\leq\ell+K$ we define $l_{i}=\ell$.
Example (continued): In the example with $K=4$ auctions from Figure 2 , $u_{\ell}=\ell+K=11$, and
$(l_{1},l_{2},l_{3},l_{4},l_{5},l_{6},l_{7},l_{8},l_{9},l_{10},l_{11})=(0,0,1,1,2,3,3,4,5,6,7).$
Using the above notation, it follows from (3.1) that
$\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))={}$ $\ln(C^{\star})+\big(\ell+\left|K_{s}\right|\big)\ln(\lambda)+\ell\ln(2)+\sum_{k\in K_{s}}\ln(t_{0,k})+\sum_{i=1}^{\ell+K}\left|Q_{i}\right|\ln(\theta_{i})$
$-\lambda\sum_{i=1}^{\ell+K}\widetilde{t}_{i}\bigg(\prod_{j=1}^{i}\theta_{j}\bigg)+1_{\{\ell>0\}}\bigg[\sum_{l=1}^{\ell}\ln(1-\theta_{u_{l}})+\sum_{i=1}^{\ell+K}(\ell-l_{i})\ln\theta_{i}\bigg],$ (3.3)
where
[…]
Note that
[…]
To maximize $\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))$, we pursue the coordinate-wise ascent approach where each iteration of the algorithm cycles through maximization of $\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))$ with respect to the co-ordinate $\theta_{i}$ (with other entries of $\boldsymbol{\theta}$ fixed at their current values) for every $1\leq i\leq\ell+K$. We now show that each of these $\ell+K$ coordinate-wise maximizers are available in closed form.
3.1 Coordinate-wise maximizers for $Lik_{A}(\lambda,\boldsymbol{\theta})$
We start with the maximization with respect to $\lambda$. Simple calculation shows that for given $F$ (or $\boldsymbol{\theta}$), $Lik_{A}\left(\lambda,\boldsymbol{\theta}\right)$ is maximized at
$\lambda=\frac{\ell+\left|K_{s}\right|}{\sum_{i=1}^{\ell+K}\widetilde{t}_{i}\bigg(\prod_{j=1}^{i}\theta_{j}\bigg)}.$
Based on the algebraic structure of $Lik_{A}(\lambda,\boldsymbol{\theta})$, we divide the coordinate-wise maximization steps into three groups: One with $\theta_{k}$ when $k\in\bf{u}$, where $\bf{u}$ is defined to be the set $\{u_{1},u_{2},\ldots,u_{\ell}\}$, the second with $\theta_{k}$ when $k\notin\bf{u}$ and $k\leq\max(\bar{S})$, and the third with $\theta_{k}$ when $k\notin u$ and $k>\max(\bar{S})$. We discuss each case in detail separately below.
Case I: Maximization w.r.t. $\theta_{i}$ for $i\in\bf{u}$. If $\bf u$} is non-empty, then $\ell>0$. For any $i\in\bf{u}$, taking derivative of the expression for $\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))$ in (3.3) w.r.t. $\theta_{i}$ and equating it to zero gives us the following
$\dfrac{\partial\big[\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))\big]}{\partial\theta_{i}}=0$
$\Leftrightarrow{}$ $-\lambda\sum_{\widetilde{i}=i}^{\ell+K}\widetilde{t}_{\widetilde{i}}\Bigg(\prod_{\underset{j\neq i}{j=1}}^{\widetilde{i}}\theta_{j}\Bigg)+\dfrac{\big(\left|Q_{i}\right|+(\ell-l_{i})\big)}{\theta_{i}}-\dfrac{1}{1-\theta_{i}}=0$
$\Leftrightarrow{}$ $-A_{i}+\dfrac{B_{i}}{\theta_{i}}-\dfrac{1}{1-\theta_{i}}=0,$ (3.4)
where
$A_{i}={}$ $\lambda\sum_{\widetilde{i}=i}^{\ell+K}\widetilde{t}_{\widetilde{i}}\Bigg(\prod_{\underset{j\neq i}{j=1}}^{\widetilde{i}}\theta_{j}\Bigg)>0$
$B_{i}={}$ $\left|Q_{i}\right|+(\ell-l_{i})>0.$ (3.5)
Since $\theta_{i}\leq 1$ and $\theta_{i}>0$, it follows that
$\dfrac{\partial\big[\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))\big]}{\partial\theta_{i}}=0$
$\implies{}$ $A_{i}\theta_{i}^{2}-\big(A_{i}+B_{i}+1\big)\theta_{i}+B_{i}=0.$ (3.6)
Since $B_{i}>0$, it follows that the discriminant of the quadratic equation (3.6), denoted by $D_{i}$, satisfies
$D_{i}={}$ $\big(A_{i}+B_{i}+1\big)^{2}-4A_{i}B_{i}$
$={}$ $\big(A_{i}-B_{i}+1\big)^{2}+4B_{i}\ >\ 0.$ (3.7)
Hence, the quadratic equation (3.6) has two real roots, namely,
$\theta_{i}=\dfrac{\big(A_{i}+B_{i}+1\big)\pm\sqrt{D_{i}}}{2A_{i}}.$ (3.8)
Since $\sqrt{D_{i}}>A_{i}-B_{i}+1$ by (3.7), it follows that
$\dfrac{\big(A_{i}+B_{i}+1\big)+\sqrt{D_{i}}}{2A_{i}}\geq\dfrac{2\big(A_{i}+1\big)}{2A_{i}}>1,$
since $A_{i}>0$. Hence the larger root with the positive sign for the square-root discriminant always lies outside the set of allowable values for $\theta_{i}$. The smaller root with the negative sign can be shown to be strictly positive since $(A_{i}+B_{i}+1)^{2}-D_{i}=4A_{i}B_{i}>0$. Also, if $A_{i}\geq B_{i}+1$, then
$A_{i}+B_{i}+1-\sqrt{D_{i}}<A_{i}+B_{i}+1\leq 2A_{i}.$
If $A_{i}<B_{i}+1$, then using $A_{i}>0$ we get
$(B_{i}+1-A_{i})^{2}=(B_{i}+1+A_{i})^{2}-4A_{i}B_{i}-4A_{i}<D_{i}$
$\Rightarrow$ $(B_{i}+1+A_{i})-\sqrt{D_{i}}<2A_{i}.$
It follows that the smaller root lies in $(0,1)$. Since
$\dfrac{\partial^{2}\big[\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))\big]}{\partial\theta_{i}^{2}}=-\dfrac{B_{i}}{\theta_{i}^{2}}-\dfrac{1}{(1-\theta_{i})^{2}}\ <\ 0,$
it follows that the smaller root is the unique maximizer of $\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))$ with respect to $\theta_{i}$. To conclude, the unique maximizer of $\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))$ with respect to $\theta_{i}$ is given by
$\widehat{\theta}_{i}=\dfrac{\big(A_{i}+B_{i}+1\big)-\sqrt{D_{i}}}{2A_{i}},$ (3.9)
where $A_{i}$ and $B_{i}$ are as defined in (3.5).
Case II: Maximization w.r.t. $\theta_{i}$ for $i\notin\bf{u}$ and $i\leq\max(\bar{S})$. For any $\theta_{i}$ with $i\notin\bf{u}$ and $i<u_{\ell}$ (other than $\theta_{1}$, which is set to $1$), the coefficient of $\ln(\theta_{i})$, given by $\left|Q_{i}\right|+1_{\{\ell>0\}}(\ell-l_{i})$, is strictly positive, while $i<u_{\ell}\leq\max(\bar{S})$. Even when $i\geq u_{\ell}$, $(\ell-l_{i})=0$, since $i\leq\max(\bar{S})$, $|Q_{i}|>0$, still the coefficient is strictly positive. Again taking derivative of the log-likelihood expression in (3.3) w.r.t. $\theta_{i}$ and equating it to zero gives us
$\dfrac{\partial\big[\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))\big]}{\partial\theta_{i}}=0$
$\Leftrightarrow{}$ $-\lambda\sum_{\widetilde{i}=i}^{\ell+K}\widetilde{t}_{\widetilde{i}}\Bigg(\prod_{\underset{j\neq i}{j=1}}^{\widetilde{i}}\theta_{j}\Bigg)+\dfrac{\big(\left|Q_{i}\right|+1_{\{\ell>0\}}(\ell-l_{i})\big)}{\theta_{i}}=0$
$\Leftrightarrow{}$ $\theta_{i}=\dfrac{\big(\left|Q_{i}\right|+1_{\{\ell>0\}}(\ell-l_{i})\big)}{A_{i}},$
where $A_{i}$ is as defined in (3.5). Note that ${(\left|Q_{i}\right|+1_{\{\ell>0\}}(\ell-l_{i}))}/{A_{i}}$ is positive but not guaranteed to be less than or equal to $1$. However, since
$\dfrac{\partial^{2}\big[\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))\big]}{\partial\theta_{i}^{2}}=-\dfrac{\big(\left|Q_{i}\right|+1_{\{\ell>0\}}(\ell-l_{i})\big)}{\theta_{i}^{2}}\ <\ 0,$
it follows that ${\partial[\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))]}/{\partial\theta_{i}}>0$, i.e., $\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))$ is an increasing function of $\theta_{i}$ for $\theta_{i}<{(\left|Q_{i}\right|+1_{\{\ell>0\}}(\ell-l_{i}))}/{A_{i}}$. Hence, the unique maximizer of $\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))$ with respect to $\theta_{i}$ is given by
$\widehat{\theta}_{i}=\min\bigg\{1,\dfrac{\big(\left|Q_{i}\right|+1_{\{\ell>0\}}(\ell-l_{i})\big)}{A_{i}}\bigg\}.$ (3.10)
Case III: Maximization w.r.t. $\theta_{i}$ for $i\notin\bf{u}$ and $i>\max(\bar{S})$. In this case $|Q_{i}|=0$ and $l_{i}=\ell$. It follows from (3.3) that $\ln(Lik_{A}(\lambda,\boldsymbol{\theta}))$ is maximized with respect to $\theta_{i}$ at 0. Hence, we set
$\widehat{\theta}_{i}=0.$ (3.11)
This amounts to estimating $F(z_{i})$ for $i>\max(\bar{S})$ by $1$. Note that any such $z_{i}$ corresponds to a starting price which is greater than the largest final selling price for all auctions in the data (including auctions where the item is sold at the starting price). Since the data offers no evidence that the support of the true valuation distribution $F$ extends up to $z_{i}$, setting the estimate of $F(z_{i})$ to $1$ indeed seems a sensible choice in this non-parametric setting.
A crucial aspect of coordinate-wise maximization of non-convex functions is effective initialization of parameter values. We first derive a generalized method of moments based initial estimator for $\lambda$, and then use it to obtain principled initial estimators for the components of $\boldsymbol{\theta}$.
3.2 Initial estimator of $\lambda$: Generalized method of moments
Consider first a single second price auction with starting price $r$, and recall that $M$ denotes the number of times the selling price changes throughout the course of the auction. Our goal is to find a function $h$ such that $E[h(M)]=\lambda$. To this end, we consider the process of consumers accessing the auction whose bid value is greater than or equal to $r$. Since we are assuming that the consumers bid their true private value, it follows that the proportion of such consumers in the population of all customers is 1F(r)1-F(r), and this “thinned” process of arriving consumers with bid values greater than $r$ is a Poisson process with rate $\lambda(1-F(r))$. Let $N_{r}$ represent the total number of consumers who access the auction in the period $0,\tau]$ and have bid values greater than the starting price $r$. Then $N_{r}\sim\mbox{Poisson}(\lambda\tau(1-F(r)))$. Moreover, given $N_{r}=n$, let $A_{i}$ ($i=1,2,\ldots,n$) represent the event that the current selling price changes after the $i^{\text{th}}$ consumer (with bid greater than $r$) accesses the auction. Let, $\textbf{1}_{A_{i}}$ be the indicator function of the occurrence of the event $A_{i}$.
Note that $E[M\mid N_{r}=0]=0=E[M\mid N_{r}=1]$, and for $n\geq 2$, we have
$E\big[M\mid N_{r}=n\big]=E\big[\text{Number of selling price changes}\mid N_{r}=n\big]={}$ $E\bigg[\sum_{i=1}^{n}\textbf{1}_{A_{i}}\Big|N_{r}=n\bigg]$
$\stackrel{{\scriptstyle(a)}}{{=}}$ $\sum_{i=2}^{n}P\big(A_{i}\mid N_{r}=n\big)$
$\stackrel{{\scriptstyle(b)}}{{=}}$ $\sum_{i=2}^{n}\frac{2(i-1)}{i(i-1)}$
$={}$ $2\sum_{i=2}^{n}i^{-1}.$ (3.12)
Here $(a)$ follows from the fact that two bids above $r$ are needed for the first change in the standing/selling price. For $(b)$, note that the arrival of $i^{\text{th}}$ consumer with bid greater than $r$ changes the selling price if and only if the corresponding bid is the highest or second highest among the $i$ starting price exceeding bids. Note that these bid values are i.i.d. with distribution $F$ truncated above $r$. There are i(i1)i(i-1) possible choices for the joint positions of the highest and second highest bids. The $i^{th}$ bid is the highest bid for (i1)(i-1) of these choices, and the second highest bid for another (i1)(i-1) choices, leading us to $(b)$. It follows from (3.12) that
$E[M]=2E\bigg[1_{\{N_{r}>1\}}\sum_{i=2}^{N_{r}}i^{-1}\bigg]=:g\big(\lambda\tau\big(1-F(r)\big)\big)\mbox{ (say). }$
Note that $1_{\{N_{r}>1\}}\sum_{i=2}^{N_{r}}i^{-1}$ is increasing in $N_{r}$, and $N_{r}$ is stochastically increasing in terms of its mean parameter $\lambda\tau(1-F(r))$. Hence $g$ is a strictly increasing function, and
$\lambda=g^{-1}(E[M])/\tau(1-F(r)).$ (3.13)
If the starting price $r$ is negligible, for example compared to the smallest final selling price seen in the data set, then it is reasonable to assume that $F(r)\approx 0$ (According to Yu et al. (2006) this is the optimal strategy for seller). Suppose now, we consider the data from $K$ independent second price auctions of identical copies of an item, with $M_{k}$ denoting the number of standing/selling price changes throughout the course of the $k^{th}$ auction, and with $r_{k}$ the starting price for the $k^{th}$ auction for $1\leq k\leq K$. Let $K_{r}$ be the collection of all auction indices with negligible starting prices. Then, it follows that from (3.13) that
$\widehat{\lambda}:=\tau^{-1}\ g^{-1}\bigg(|K_{r}|^{-1}\sum_{k\in K_{r}}M_{k}\bigg),$
should be a reasonable generalized method of moments estimator for $\lambda$.
Of course, the function $g$ is not available in closed form and needs to be computed using numerical methods. A natural approach, given the definition of $g$ as a Poisson expectation, is Monte Carlo. Indeed, we computed $g(x)$ for every $x$ on a fine grid (with spacing $0.1$) ranging from 0 to $5$. This Monte Carlo computation of $g$ is a one-time process that required minimal computational effort. The resulting plot of $g$ is provided in Figure 3.
Figure 3: Plot of $g$ over the interval $\left[0,5\right]$.
3.3 Constructing an initial estimate $\widehat{F}_{init}$ of $F$ (and $\boldsymbol{\theta}^{(0)}$ of $\boldsymbol{\theta}$) based exclusively on final selling prices and first observed bids
The details of all the steps of the coordinate ascent maximization algorithm for $Lik_{A}$ are explicitly derived above in Section 3.1. However, as mentioned previously, a crucial detail which needs to be worked out is a ‘good’ initial starting point for the algorithm. Especially for highly non-convex maximization such as in the current setting, the choice of the initial/starting value can play a critical role in the quality of the final estimate produced by the coordinate ascent algorithm. In this section, we construct an initial estimate of $F$ based on the empirical distribution functions of both the final selling prices and the first observed bids (i.e., the price when for the first time the standing price jumps to a higher value from its respective starting price), respectively. Note that the methodology developed in George and Hui (2012) also relies exclusively on the final selling prices. However, that approach requires the knowledge of the total number of consumers accessing each of the $K$ auctions. We do not assume this knowledge in the current setting, and need to overcome this additional challenge. Also, as stated above, we will use the first non-starting standing prices (first observed bids) to improve the quality of our initial estimator of $F$.
Once the initial estimate $\widehat{F}_{init}$ of $F$ is constructed, we can easily construct the initial estimate $\boldsymbol{\theta}^{(0)}:=(\theta_{1}^{(0)},\theta_{2}^{(0)},\ldots,\theta_{\ell+K}^{(0)})^{T}$ of $\boldsymbol{\theta}$ using (2.3) as follows:
$\theta_{i}^{(0)}=\frac{1-\widehat{F}_{init}(z_{i})}{1-\widehat{F}_{init}(z_{i-1})},\ \forall\ 1\leq i\leq(\ell+K).$ (3.14)
We now describe in detail the various steps involved in construction of $\widehat{F}_{init}$.
Step I: Construct an estimate of $F$ based on the empirical distribution function of only the final selling prices of auctions with relatively small starting prices. First, consider a single second price auction with starting price $X_{0}=r$. As in Section 3.2 consider the process of consumers accessing the auction whose bid value is greater than or equal to $r$, and let $N_{r}$ represents the number of such consumers that access the auction in the period $0,\tau]$. As observed in Section 3.2, this thinned process of arriving consumers is a Poisson process with rate $\lambda(1-F(r))$, and $N_{r}\sim Poisson(\lambda\tau(1-F(r)))$. We derive the conditional cdf of the final selling price $X_{M}$ given $X_{0}=r,N_{r}\geq 2$ as a function of $F(x)$. For this purpose, note that
$P(X_{M}\leq x\mid X_{0}=r,N_{r}\geq 2)={}$ $\sum_{n=2}^{\infty}P(X_{M}\leq x\mid N_{r}=n,X_{0}=r)P(N_{r}=n\mid N_{r}\geq 2)$
$={}$ $\sum_{n=2}^{\infty}P(X_{M}\leq x,R>x\mid X_{0}=r,N_{r}=n)P(N_{r}=n\mid N_{r}\geq 2)$
$+\sum_{n=2}^{\infty}P(X_{M}\leq x,R\leq x\mid X_{0}=r,N_{r}=n)P(N_{r}=n\mid N_{r}\geq 2).$ (3.15)
Recall that $R$ denotes the maximum placed bid during the course of the auction (we do not observe it), and that the valuation distribution of customers arriving in the thinned Poisson process discussed above is the truncated version of $F$ at $r$, denoted by
$F_{r}(x):=\frac{F(x)-F(r)}{1-F(r)}\ 1_{\{x>r\}}.$
Now, note that the event $\{X_{M}\leq x,R>x\}$ is equivalent to the constraint that the largest order statistic of the valuations of all the customers arriving via the thinned Poisson process is greater than $x$, but second largest order statistic is less than or equal to $x$. Similarly, the event $\{X_{M}\leq x,R\leq x\}$ is equivalent to the constraint that the largest order statistic of the valuations of all the customers arriving via the thinned Poisson process is less than or equal to $x$. With $\lambda_{r}=\lambda(1-F(r))$, it follows from (3.15) that
$P(X_{M}\leq x\mid X_{0}=r,N_{r}\geq 2)$
$={}$ $\frac{1}{P(N\geq 2)}\bigg[\sum_{n=2}^{\infty}\frac{n(F_{r}(x))^{n-1}(1-F_{r}(x))\exp(-{\lambda_{r}}\tau)({\lambda_{r}}\tau)^{n}}{n!}+\sum_{n=2}^{\infty}\frac{(F_{r}(x))^{n}\exp(-{\lambda_{r}}\tau)({\lambda_{r}}\tau)^{n}}{n!}\bigg]$
$={}$ $\frac{1}{P(N\geq 2)}\bigg[{\lambda_{r}}\tau\exp(-{\lambda_{r}}\tau)(1-F_{r}(x))\sum_{n=2}^{\infty}\frac{({\lambda_{r}}\tau F_{r}(x))^{n-1}}{(n-1)!}+\exp(-{\lambda_{r}}\tau)\sum_{n=2}^{\infty}\frac{({\lambda_{r}}\tau F_{r}(x))^{n}}{n!}\bigg]$
$={}$ $\frac{{\lambda_{r}}\tau\exp(-{\lambda_{r}}\tau)(1-F_{r}(x))\big(\exp({\lambda_{r}}\tau F_{r}(x))-1\big)+\exp(-{\lambda_{r}}\tau)\big(\exp({\lambda_{r}}\tau F_{r}(x))-{\lambda_{r}}\tau F_{r}(x)-1\big)}{P(N\geq 2)}$
$={}$ $\frac{\exp(-{\lambda_{r}}\tau)\ \Big({\lambda_{r}}\tau(1-\eta)\big(\exp({\lambda_{r}}\tau\eta)-1\big)+\exp({\lambda_{r}}\tau\eta)-{\lambda_{r}}\tau\eta-1\Big)}{1-\exp(-{\lambda_{r}}\tau)-{\lambda_{r}}\tau\exp(-{\lambda_{r}}\tau)},$
where $\eta=F_{r}(x)$. Let,
$G_{{\lambda_{r}}}(\eta):=\frac{\exp(-{\lambda_{r}}\tau)\ \Big({\lambda_{r}}\tau(1-\eta)\big(\exp({\lambda_{r}}\tau\eta)-1\big)+\exp({\lambda_{r}}\tau\eta)-{\lambda_{r}}\tau\eta-1\Big)}{1-\exp(-{\lambda_{r}}\tau)-{\lambda_{r}}\tau\exp(-{\lambda_{r}}\tau)}.$ (3.16)
Note that
$\frac{d}{d\eta}G_{{\lambda_{r}}}(\eta)={}$ $\frac{({\lambda_{r}}\tau)^{2}(1-\eta)\exp({\lambda_{r}}\tau\eta)-{\lambda_{r}}\tau\big(\exp({\lambda_{r}}\tau\eta)-1\big)+{\lambda_{r}}\tau\exp({\lambda_{r}}\tau\eta)-{\lambda_{r}}\tau}{\exp({\lambda_{r}}\tau)-(1+{\lambda_{r}}\tau)}$
$={}$ $\frac{({\lambda_{r}}\tau)^{2}(1-\eta)\exp({\lambda_{r}}\tau\eta)}{\exp({\lambda_{r}}\tau)-(1+{\lambda_{r}}\tau)}\ >\ 0\quad\text{for}\ \eta\in(0,1).$ (3.17)
It follows that $G_{\lambda_{r}}$ is a strictly increasing function for $\eta\in[0,1]$.
Now, coming back to our setting with $K$ independent auctions, suppose that we have multiple auctions with a given starting price $r$ (or close to $r$) where the item is sold above the starting price. Then based on the Glivenko-Cantelli lemma, (3.16) and (3.17), we can use the function $G_{\lambda_{r}}^{-1}$ (with an appropriate estimate of $\lambda_{r}$) applied to empirical cdf of the final selling prices of these auctions to estimate $F_{r}(x)$ for $x>r$. Setting the estimate of $F(r)$ to be zero, we can then obtain an estimate of $F(x)$ for $x>r$. Clearly, we would like to choose $r$ to be as small as possible.
With this background, let $r_{q}$ denote the $q^{th}$ quantile of the starting prices among $\{r_{1},r_{2},\cdots,r_{K}\}$. Here $q\in(0,1)$ is a user-specified constant, and we denote by $V(q)$ the set of indices of starting prices which lie within $0,r_{q}]$. Ideally, one would like to have a reasonable number of auctions with very small/negligible starting prices. For example, in the Xbox data analyzed in Section 5, roughly $25\%$ of the auctions have a starting price less than $\mathdollar 1$ (the smallest final selling price is $\mathdollar 28$). Let
$G_{SP}(x):=\frac{1}{|V(q)|}\sum_{k\in V(q)}1_{\{X_{m_{k},k}\ \leq\ x\}},$
be the empirical distribution function of the final selling prices for auctions in $V(q)$. Based on the above discussion we construct the estimator $\widehat{F}_{SP}$ of $F$ as
$\widehat{F}_{SP}(x)=G_{\widehat{\lambda}}^{-1}(G_{SP}(x)),\ \forall\ x\in\mathbb{R}.$ (3.18)
Here $\widehat{\lambda}$ is our initial estimate of $\lambda$. In fact, $\widehat{F}_{SP}(x)=0$ for all values below the smallest final selling price for auctions corresponding to $V(q)$. There are likely many observed standing prices in the $K$ auctions which are below this smallest final selling price, and these values can/should be used to improve the estimator $\widehat{F}_{SP}$. This process is described in the next step.
Step II: Incorporate the first non-starting standing prices into the construction of the initial estimate of $F$. Consider again, to begin with, a single second price auction with starting price $r$, and the associated thinned Poisson process of arriving consumers with valuation greater than $r$. Letting $Y_{1},Y_{2}$ represent valuations of the first two arriving consumers in the thinned process, we have
$P(X_{1}\leq x\mid X_{0}=r,N_{r}\geq 2)=P(\min\{Y_{1},Y_{2}\}\leq x)=1-(1-F_{r}(x))^{2}.$ (3.19)
Similar to Step I, let
$G_{FP}(x):=\frac{1}{|V(q)|}\sum_{k\in V(q)}1_{\{X_{1,k}\ \leq\ x\}},$
be the empirical distribution function of the first non-starting standing prices for auctions in $V(q)$. Based on (3.19), we construct the estimator $\widehat{F}_{FP}$ of $F$ as
$\widehat{F}_{FP}(x)=1-\sqrt{1-G_{FP}(x)},\ \forall\ x\in\mathbb{R}.$ (3.20)
Note that $G_{FP}(z_{1})=0$, which implies that $\widehat{F}_{FP}(z_{1})=0$, where $z_{1}$ is the smallest starting price. However, $\widehat{F}_{FP}(x)>0$ when $x$ is larger than the smallest first non-starting standing price among auctions in $V(q)$. This smallest non-starting standing price is often considerably smaller than the smallest final selling price, and hence $\widehat{F}_{FP}$ can be combined with $\widehat{F}_{SP}$ of Step I, to get a better initial estimate of $F$ as follows.
Step III: Combining the two different initial estimates, namely, $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$. Let $p_{1}$ and $p_{2}$ respectively represent the largest non-starting standing price and the smallest final selling price for auctions in $V(q)$. As discussed previously, $\widehat{F}_{SP}$ underestimates $F$ below $p_{2}$ and $\widehat{F}_{FP}$ overestimates $F$ above $p_{1}$. Let $c$ be the largest real number $\leq\min\{p_{1},p_{2}\}$ such that $\widehat{F}_{FP}(c)\leq\widehat{F}_{SP}(p_{1})$. Then, we define a function $\widehat{F}_{(0)}$ based on $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$ as follows:
$\widehat{F}_{(0)}(x)=\begin{cases}\widehat{F}_{FP}(x)&\text{if}\ \ x\leq c\\ \widehat{F}_{SP}(x)&\text{if}\ \ x>p_{2}\\ \widehat{F}_{FP}(c)+\Big(\frac{\widehat{F}_{SP}(p_{1})-\widehat{F}_{FP}(c)}{p_{1}-c}\Big)(x-c)&\text{if}\ \ c<x\leq p_{2}.\end{cases}$ (3.21)
This function $\widehat{F}_{(0)}$ in (3.21) combines the strengths of the two estimators $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$, and gives a more balanced estimator of $F$ over all regions. Finally, since $G_{FP},G_{SP}$ are step functions, so are $\widehat{F}_{FP},\widehat{F}_{SP}$. It follows based on (3.21) that $\widehat{F}_{(0)}$ is a step function as well, and has jumps only at the first non-starting standing prices and final selling prices for auctions in $V(q)$. A continuous version of this estimator, denoted by $\widehat{F}_{init}$ can be obtained by linear interpolation of the values between the jump points. Since $z_{1}\leq c$, it follows that $\widehat{F}_{(0)}(z_{1})=\widehat{F}_{FP}(z_{1})=0$.
3.4 The Coordinate ascent algorithm for maximizing $Lik_{A}$
All the developments and derivations in the earlier subsections can now be compiled and summarized via the following coordinate ascent algorithm to maximize $Lik_{A}(\lambda,\boldsymbol{\theta})$.
Algorithm 3.1.
Coordinate ascent algorithm:
  1. Step 1.
    Start with initial value $\lambda^{(0)}=\widehat{\lambda}$ (Section 3.2) and $\boldsymbol{\theta}^{(0)}=(\theta_{1}^{(0)},\theta_{2}^{(0)},\ldots,\theta_{\ell+K}^{(0)})^{T}$ (Section 3.3), and a user defined tolerance level $\epsilon>0$. Note $\theta_{1}^{(0)}=1$ since $\widehat{F}_{(0)}(z_{1})=0$.
  2. Step 2.
    Set $m=0$.
  3. Step 3.
    Set $\theta_{1}^{(m+1)}=1$ (for identifiability). For any $2\leq i\leq(\ell+K)$, sequentially obtain $\theta_{i}^{(m+1)}$ from (3.9), (3.10) and (3.11) by using the coordinate values in
    $(\theta_{1}^{(m+1)},\ldots,\theta_{i-1}^{(m+1)},\theta_{i+1}^{(m)},\ldots,\theta_{\ell+K}^{(m)})^{T}$ to compute $A_{i},B_{i},C_{i}$.
  4. Step 4.
    Update $\lambda^{(m+1)}=\big(\ell+\left|K_{s}\right|\big)/\left(\sum_{i=1}^{\ell+K}\widetilde{t}_{i}\bigg(\prod_{j=1}^{i}\theta_{j}^{(m+1)}\bigg)\right)$
  5. Step 5.
    If
    $\ln\big(Lik_{A}\big(\lambda^{(m+1)},\boldsymbol{\theta}^{(m+1)}\big)\big)-\ln\big(Lik_{A}\big(\lambda^{(m)},\boldsymbol{\theta}^{(m)}\big)\big)>\epsilon,$
    set $m\leftarrow m+1$. Otherwise, go to Step6~6.
  6. Step 6.
    Set $\widehat{\boldsymbol{\theta}}_{MLE}=\boldsymbol{\theta}^{(m+1)}$ and $\widehat{\lambda}_{MLE}=\lambda^{(m+1)}$.
Once we get $\widehat{\boldsymbol{\theta}}_{MLE}$, we can easily get the corresponding maximum likelihood estimator of $F$ as follows
$\widehat{F}_{MLE}(z_{i})=1-\prod_{j=1}^{i}\widehat{\theta}_{MLE,j}\ ,\ \forall\ 1\leq i\leq(\ell+K).$ (3.22)
As explained above, the adapted objective function $Lik_{A}$ and the search space $\mathcal{F}_{\bf z}$ of CDFs with relevant jump discontinuities are artifacts of the semi-parametric approach that we pursue. However, once the estimates of the valuation distribution at elements of $\bf z$} are obtained using Algorithm 3.1, a continuous estimate on the entire valuation distribution can be constructed via interpolation. In particular, we use the values of $\widehat{F}_{(MLE)}$ at $z_{i}$’s, $\widehat{F}_{(MLE)}(0)=0$, and linear interpolation to construct a continuous estimator of the population valuation distribution over the entire real line.
4 Simulation study
In this section we consider various choices of the true underlying valuation distribution $F$, e.g., uniform, piecewise uniform, pareto, gamma, and beta distributions, which are commonly used in marketing research. We then illustrate and compare the performance of the MLE $\widehat{F}_{MLE}$, the initial estimate $\widehat{F}_{init}$ and the Pólya Tree estimate $\widehat{F}_{PT}$ estimate from George and Hui (2012) with the corresponding true valuation distribution $F$. Note that the Bayesian methodology in George and Hui (2012) (which uses only the final selling prices in each auction) requires the knowledge of the total number of consumers accessing the auction. This does not hold in our motivating application, but in order to compare these methods we will also assume such a knowledge in our synthetic data evaluations for $\widehat{F}_{PT}$ below.
4.1 Data generation
We conducted five sets of simulation experiments, each using data simulated from a different choice of the underlying valuation distribution $F$. The cumulative distribution functions corresponding to the five choices of true underlying valuation distributions are shown in Figure 4.
Figure 4: “True” underlying valuation distribution functions used in the simulation studies.
[…]
From each of the five true underlying $F$’s, we consider two settings, with $K=100$ and $K=1000$ independent auctions of identical copies of an item. Varying the sample size here sheds light on the relationship between the sample size and the precision of the MLE $\widehat{F}_{MLE}$, the initial estimate $\widehat{F}_{init}$ and the Pólya Tree estimate $\widehat{F}_{PT}$ in estimating the true valuation distribution $F$. For each auction, we took the auction window ($\tau$) to be 100 units, and the constant rate ($\lambda$) of arrival of bidders to be equal to $1$. We then simulated the inter-arrival times between bidders from an exponential distribution with rate parameter $\lambda=1$, and drew the bidders’ valuations from $F$, keeping track of the entire sequence of standing prices and the intermediate times between jumps in the standing price for all independent auctions involved. This sequence of standing prices throughout the course of the auction, and the intermediate times between standing price changes are then treated as the observed dataset that is subsequently used to compute the initial estimator and the MLE. For each choice of true $F$ and number of auctions $K$, $100$ replicated datasets are generated.
Since the data is generated to be consistent with the modeling assumptions, one expects $\widehat{F}_{MLE}$, which utilizes all available information, to have a superior performance than the initial estimator $\widehat{F}_{init}$, which only uses the final selling price and first non-starting standing price for each auction, and the Pólya Tree estimator $\widehat{F}_{PT}$ of George and Hui (2012), which only uses final selling prices. The goal of these simulations is to examine extensively how much improvement can be obtained from our proposed method by incorporating the additional information in a variety of settings.
4.2 Simulation Results
For each replicated dataset generated (as described in the previous subsection), we apply our semi-parametric methodology to obtain the initial estimate $\widehat{F}_{init}$, the MLE $\widehat{F}_{MLE}$, and the Pólya Tree estimator $\widehat{F}_{PT}$ in George and Hui (2012). The goal is to compare the accuracy of each of these estimators with respect to the respective true valuation distribution $F$.
[…]
Figure 5: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be piece-wise Uniform) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
Figure 6: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be Uniform(1,20)(1,20)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
Figure 7: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be Pareto(location = 3,dispersion = 100)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
Figure 8: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be Gamma(shape = 10, rate = 2)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
Figure 9: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be Beta(2,2)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
The above plots based on single chosen replicates are illustrative, but need to be complemented with performance evaluation averaged over all the $100$ replicates in each of the $10$ simulation settings. In Table 2, for each simulation setting, we provide both the Kolmogorov-Smirnov distance (KS-distance) and the Total variation distance (TV-distance) between the true valuation distribution $F$ and various estimators. In this comparison, we also include two parametric methods, the Gamma-F method and the Truncated Normal-F (TN-F) method, used in George and Hui (2012). Both methods use final selling prices, and similar to Pólya Tree, require the knowledge of the total number of potential bidders who access the auction. The distances for all five methods (MLE, Initial, Pólya Tree, Gamma-F and TN-F) are averaged over the $100$ respective replications. The results show that the MLE (based on the entire collection of standing prices) uniformly outperforms the initial estimator (based only on final selling prices and first non-starting standing prices) and the other three final selling price based estimators (Pólya Tree, Gamma-F and TN-F) in all the simulation settings. This strongly suggests that if the additional assumptions of single bidding and constant arrival rate seem to largely hold, it is worth using the proposed methodology which incorporates the extra information available in the form of all standing prices within the auction period.
KS distance
Distribution $K$     MLE Initial Pólya Tree Gamma-F TN-F
Uniform 100 0.048 0.138 0.050 0.266 0.732
1000 0.015 0.051 0.050 0.302 0.727
Piecewise-Uniform 100 0.048 0.104 0.239 0.385 0.708
1000 0.015 0.035 0.237 0.397 0.716
Pareto 100 0.048 0.114 0.280 0.061 0.131
1000 0.017 0.041 0.251 0.021 0.124
Gamma 100 0.045 0.134 0.290 0.314 0.239
1000 0.014 0.051 0.296 0.348 0.229
Beta 100 0.054 0.146 0.121 0.281 0.488
1000 0.018 0.057 0.123 0.382 0.478
Total Variation
Uniform 100 0.075 0.211 0.058 0.296 0.738
1000 0.033 0.089 0.060 0.313 0.734
Piecewise-Uniform 100 0.091 0.151 0.486 0.445 0.674
1000 0.065 0.082 0.479 0.433 0.691
Pareto 100 0.057 0.174 0.281 0.057 0.13
1000 0.019 0.069 0.249 0.021 0.123
Gamma 100 0.067 0.228 0.317 0.316 0.241
1000 0.022 0.083 0.309 0.351 0.229
Beta 100 0.074 0.237 0.133 0.288 0.483
1000 0.025 0.090 0.129 0.385 0.472
Table 2: Kolmogorov-Smirnov (KS) distance and Total variation distance between various estimators ($\widehat{F}_{MLE}$, $\widehat{F}_{init}$, Pólya Tree estimator(PT), the Gamma method based on final selling prices - Gamma-F, the Truncated-Normal method based on final selling prices - TN-F) and the true valuation distribution $F$, averaged over 100 replications within each of the 10 simulation settings.
For uncertainty quantification, we provide the coverage rate for 90% HulC-based intervals for the proposed MLE and Init approaches, and also the coverage rate for 90% credible intervals for the Pólya Tree (PT) approach of George and Hui (2012), under varied simulation settings. Note that for every approach, we compute the coverage rate as the proportion of replications for which the entire true distribution function is contained within the respective upper and lower confidence bands (with Bonferroni correction used to adjust for multiple testing). The results show that the proposed MLE approach provides significantly better coverage than the other two approaches. We also computed the band area (area between the lower and upper confidence bands) for each of the approaches, and the respective values (averaged over all replications) are provided in Table 3 as well. These values reveal that the Init approach yields significantly wider confidence bands than the MLE approach, yet suffers from lower coverage. On the other hand, the PT approach yields significantly narrower bands compared to the MLE approach, but has zero empirical coverage. So overall, the proposed MLE approach strikes a much better balance overall than the other two estimators.
The HulC approach assumes median unbiasedness of the underlying estimators, but $\widehat{F}_{init}$ and $\widehat{F}_{MLE}$ are not median unbiased - adjusting for median bias in HulC can be computationally expensive and leads to wider confidence bands. Note that near the boundaries, where the value of the underlying true CDF is very close to zero or one, median unbiasedness is a tall ask for these non-parametric/semi-parametric estimators. Initial studies did show that the estimated median bias for $\widehat{F}_{MLE}$ was very small across the simulation settings for points away from the boundary (points within 5th to 95th percentile range). The last three columns of Table 3 provide ‘truncated coverage’ for the three methods across all settings, i.e., coverage calculation is restricted only to points within the 5th to 95th percentile range of the true distribution $F$. The results show that the observed coverage for $\widehat{F}_{init}$ and $\widehat{F}_{MLE}$ moves closer to the nominal coverage of 90% in many settings when we restrict away from the boundary. The comparative performance of the three methods remains the same. These results suggest that is not much advantage, in terms of coverage, by adjusting for median bias of the proposed MLE estimator in HulC.
Coverage Band Area Truncated Coverage
Distribution $K$ MLE Initial PT MLE Initial PT MLE Initial PT
Uniform 100 0.65 0.63 0.00 3.61 6.66 0.45 0.92 0.90 0.00
1000 0.94 0.84 0.00 1.09 3.14 0.42 0.94 0.84 0.00
Piecewise- 100 0.97 0.83 0.00 0.61 0.98 0.09 0.90 0.77 0.00
Uniform 1000 0.98 0.96 0.00 0.18 0.38 0.09 0.98 0.95 0.00
Pareto 100 0.56 0.01 0.00 1.25 2.90 0.27 0.89 0.01 0.00
1000 1.00 0.93 0.00 0.45 1.18 0.20 0.97 0.88 0.00
Gamma 100 0.42 0.05 0.00 0.90 2.09 0.20 0.89 0.07 0.00
1000 0.97 0.81 0.00 0.30 0.92 0.18 0.97 0.86 0.00
Beta 100 0.36 0.18 0.00 0.15 0.32 0.02 0.96 0.65 0.00
1000 0.91 0.74 0.00 0.05 0.14 0.02 0.97 0.81 0.00
Table 3: Empirical coverage, average band area, and empirical truncated coverage for the HulC confidence intervals of estimators $\widehat{F}_{MLE}$, $\widehat{F}_{init}$, and for the credible interval of estimator $\widehat{F}_{PT}$ (denoted as PT) , averaged over 100 replications within each of the 10 simulation settings. Truncated coverage refers to coverage calculated only on points within the 5th to 95th percentile range of the true distribution $F$.
4.3 Evaluation based on profit-maximizing price
As noted in George and Hui (2012), a key advantage of knowing the valuation distribution $F$ is that it enables managers to determine the profit-maximizing price for a product in a given market. In particular, if $c$ is the production cost, then the expected (per person) profit at a price $x$ is given by
$\pi(x)=(1-F(x))(x-c)$
In particular, $\pi(x)$ is the total expected profit divided by the total number of potential customers in the underlying population. As George and Hui (2012) note, $\pi(x)$ is the product of the proportion of potential customers with valuation above price $x$ (i.e., 1F(x)1-F(x)), and the profit for each sale (i.e., xcx-c). Hence, the profit maximizing price $x_{F}$, is defined as
$x_{F}=\mbox{argmax}_{x}(1-F(x))(x-c).$
Based on an estimate $\widehat{F}$ of $F$, the underlying profit maximizing price $x_{F}$ can be estimated by
$x_{\widehat{F}}=\mbox{argmax}_{x}(1-\widehat{F}(x))(x-c).$
Based on these considerations, for each of the ten simulation settings—corresponding to five choices of the true valuation distribution $F$ and two choices of the total number of auctions KK—we now additionally evaluate each method along two dimensions: (a) the proximity of the estimated profit‑maximizing price $x_{\widehat{F}}$ to the true profit‑maximizing price $x_{F}$, and (b) the proximity of the estimated (per person) maximum profit
$\widehat{\pi}(x_{\widehat{F}})=\bigl(1-\widehat{F}(x_{\widehat{F}})\bigr)\bigl(x_{\widehat{F}}-c\bigr)$
[…]
We use five methods in our comparison. These include the proposed MLE approach, the Initial approach, the Pólya Tree approach of George and Hui (2012) and two parametric approaches based on final selling prices used in George and Hui (2012) - the Gamma-F method, and the Truncated Normal-F (TN-F) method.
Figure 10 provides box plots for the estimated (per person) profit maximizing price $x_{\widehat{F}}$ of each method in each of the 10 simulation settings. The production cost $c$ was set to the 25% quantile of the underlying true distribution, and the results described below exhibit similar patterns even when $c$ varies from the 10% to the 50% quantile (see Supplemental Section S5 for more details).
It can be seen that the estimated profit‑maximizing price produced by the MLE method is generally very close to the true optimal price, while exhibiting lower variability than the Initial method. In contrast, the Pólya tree method tends to overestimate the optimal price across nearly all settings (Piecewise Uniform being the sole exception), whereas the Gamma method typically underestimates it (Gamma being the sole exception). The Truncated-Normal method tends to have a much larger bias compared to the other methods.
[…]
In contrast, the Gamma method systematically underestimates the maximum profit in nearly all settings, with the Pareto case being the sole exception. The Pólya tree estimator displays more variable behavior: it tends to overestimate profits most severely in the Pareto setting, while underestimating profits in the Uniform and the Piecewise Uniform settings. The Truncated-Normal method performs the worst overall, showing poor accuracy relative to the other estimators in most simulation settings.
Figure 10: Box plots for the estimated profit maximizing price $x_{\widehat{F}}$ for each of the five estimators, $\widehat{F}_{MLE}$, $\widehat{F}_{init}$, Pólya tree estimator (PT), the Gamma method based on final selling prices (Gamma-F), and the Truncated-Normal method based on final selling prices (TN-F). Results are based on 100 replications within each of the 10 simulation settings. The dashed lines represent the true profit maximizing price $x_{F}$ in each scenario.
[…]
Distribution $K$     MLE Initial Pólya Tree Gamma-F TN-F
Uniform 100 +0.064 +0.235 -0.119 -1.094 +7.250
1000 +0.022 +0.115 -0.096 -1.286 +7.164
Piecewise-Uniform 100 +0.006 +0.007 -0.240 -0.157 +1.399
1000 +0.001 +0.008 -0.242 -0.192 +1.426
Pareto 100 +0.027 +0.191 +0.915 +0.033 +0.370
1000 +0.009 +0.055 +0.739 -0.006 +0.347
Gamma 100 +0.013 +0.118 +0.022 -0.256 -0.117
1000 +0.003 +0.026 -0.014 -0.277 -0.117
Beta 100 +0.002 +0.017 -0.003 -0.050 -0.071
1000 +0.000 +0.004 -0.005 -0.063 -0.071
[…]
5 Empirical application
In this section we apply our method to estimate the true valuation distribution of an Xbox based on actual data obtained from second-price auctions on eBay. In Section 5.1, we provide an overview of the data, and discuss features and adjustments to ensure its suitability for the methodology developed in the paper. In Section 5.2, we apply our semi-parametric methodology on the data set and present the findings, and perform additional performance analysis.
5.1 Data overview
The data set on eBay on online auctions of Xbox game consoles was obtained from the Modeling Online Auctions data repository. More specifically, we focus on a data set which provides information for $93$ online auctions of identical Xbox game consoles where each auction lasts for $7$ days. For each auction, a user’s bid is recorded only if it changes the standing price in the auction. For each such bid, the following information is provided: auctionid (unique auction identifier), bid (dollar value of the bid), bidtime (the time, in days, that the bid was placed), bidder (bidder eBay username), bidderrate (internal eBay rating of the bidder), openbid (the starting price for the auction, set by the seller), and price (the final selling price for the auction). While the standing price values throughout the course of the auction were not directly provided, they can be easily inferred from the successful bid values from the bid column and the starting price. Also, the bidtime column directly provides the sequence of times at which there is a change in the standing price.
[…]
5.2 Analysis of Xbox data
Using the Xbox 7-day auctions dataset with slight modifications as mentioned in Section 5.1, we now compute the initial estimate $\widehat{F}_{init}$ and the MLE $\widehat{F}_{MLE}$. For the estimation of $\lambda$ (see Section 3.2) we need to choose a subset of auctions whose starting prices are negligible in the given context. We found that the smallest final selling price in all the auctions is \$25 and the median final selling price in all the auctions is \$125. Given this, we chose all auctions with starting price less than \$1 ($16$ out of $93$) for obtaining the generalized method of moments based estimator of $\lambda$, and also for the computing the final selling price based estimator $\widehat{F}_{SP}$ (see Step I in Section 3.3). Recall that $\widehat{F}_{SP}$ is one of the components used to compute the initial estimator $\widehat{F}_{init}$.
The plots of the initial estimate $\widehat{F}_{init}$ and the MLE $\widehat{F}_{MLE}$ of the (unknown) true valuation distribution along with the corresponding $90\%$ HulC confidence regions are provided in Figure 11. Similar to the phenomenon observed in the simulations in Section 4.2, we notice that the HulC confidence region of $\widehat{F}_{MLE}$ is narrower than that of $\widehat{F}_{init}$, indicating comparatively smaller variance of $\widehat{F}_{MLE}$. Another interesting observation is that the curves for these two estimates cross exactly once, with $\widehat{F}_{MLE}(x)$ dominated by $\widehat{F}_{init}(x)$ after the crossing point, and vice-versa before the crossing point. This implies that $\widehat{F}_{init}$ stochastically dominates $\widehat{F}_{MLE}$. In other words, the final selling price/first non-starting standing price based initial estimator signifies higher Xbox valuations than the MLE estimator based on the entire collection of standing prices throughout the course of the auctions.
Unlike the simulation setting, the true valuation distribution is obviously not known here. However, we still undertake a limited performance evaluation and comparison exercise for the two approaches. As discussed previously in Section 4.1, if the modeling assumptions are largely unviolated (which seems to be the case) one would expect the MLE to do better than the initial estimator. The goal of this limited evaluation is again to understand the amount of improvement, and also to examine the stability of both estimators. For this purpose, we split the entire Xbox dataset into training and test sets. In particular, we consider two choices of splits namely, $1:1$ and $2:1$, for the ratio of auctions in training vs. test data. For each splitting proportion, 100 random splits are performed. For each split, estimates $\widehat{F}_{MLE}$ and $\widehat{F}_{init}$ are obtained from the training data, along with the estimate $\widehat{\lambda}_{MLE}$ of the Poisson arrival rate. For each test auction and each of two estimates of $F$ (init and MLE), 100 pseudo auctions with the same starting price are generated. The mean squared error of the corresponding 100 final selling prices is computed around the observed final selling price in the test auction. These relative mean squared errors are then averaged over all the test auctions. The mean squared errors are finally averaged over these 100 replications/splits to obtain an overall MSE for each of the two methods. The MSEs for both $\widehat{F}_{MLE}$ and $\widehat{F}_{init}$ are normalized by the MSE of the preliminary estimator $\widehat{F}_{FP}$ (based solely on the first non-starting standing prices). The resulting normalized MSE values are provided in Table 5. In both settings, MLE yields a better estimate of the final selling price compared to the init.
Figure 11: Plot of Initial estimate (solid blue line, based only on final selling prices and first observed bids) vs. MLE (solid red line, based on entire sequence of standing prices) of F and their corresponding $90\%$ HulC confidence regions (dotted blue and red lines) for the Xbox 7-day auctions dataset.
train : test $1:1$ $2:1$
$MSE(\widehat{F}_{init})/MSE(\widehat{F}_{FP})$ $0.1535$ $0.1565$
$MSE(\widehat{F}_{MLE})/MSE(\widehat{F}_{FP})$ $0.1265$ $0.1218$
Table 5: Mean squared error (MSE) between data generated based on starting price in the test set and estimated $\widehat{\lambda}_{MLE}$, $\widehat{\lambda}_{init}$, $\widehat{F}_{MLE,\ test}$ and each of $\widehat{F}_{init,\ training}$, $\widehat{F}_{MLE,\ training}$ and $\widehat{F}_{PT,\ training}$ (with number of bidders generated from MLE model) respectively, compared with data generated from $\widehat{F}_{FP,\ training}$, averaged over $100$ replications of the random split with same proportion, and in each replicate the data is generated 100 times based on the same set of starting price.
6 Discussion and future research
In this paper we have developed a semi-parametric methodology for estimating the consumer valuation distribution using second price auction data. Unlike the approach in George and Hui (2012), our methodology uses the collection of current selling price values throughout the course of the auctions, and does not require knowledge of the total number of bidders accessing the auction. Extensive simulations demonstrate that, when the modeling assumptions are true, using our approach can lead to significantly better performance than estimators based on just final selling prices and first observed bids. Two additional assumptions (compared to George and Hui (2012)) which preclude multiple bidding and postulate constant rate of arrival of the consumers to the auction are needed. Many real-life second price auctions see only minor departures from these assumptions, which are supported by economic theory. However, if there is evidence of major violation, results from the proposed methodology should be used cautiously. Generalizing our methodology by relaxing one or both of these assumptions is a topic of future research. One possible direction which we plan on exploring is allow two different rates for the bidder arrival process, with a transition between these two rates happening sometime during the auction period.
Another pertinent direction for future research is the incorporation of covariates in the model; see, for example, P. Bajari and A. Hortacsu (2003). The population valuation distribution for the same product might depend on factors such as condition of the item and seller statistics. One possible path for incorporating this in our model is through a semi-parametric approach where a linear combination of the relevant covariates is added as a location parameter for the cdf $F$. In particular, the population valuation cdf, evaluated at $t$, with $K$ relevant covariate values $x_{1},x_{2},\ldots,x_{K}$ is given by $F\left(t-\sum_{i=1}^{K}\beta_{i}x_{i}\right)$, see e.g., Groeneboom and Hendrickx (2018). The corresponding likelihood maximization will now involve the unknown regression parameters $\{\beta_{i}\}_{i=1}^{K}$ in addition to $\lambda$ and $F$. While the conditional maximization of $\lambda$ and $F$ will involve similar calculations as those developed in this paper, conditional maximization of the regression coefficients will need more careful thought and analysis.
{funding}
Rohit Patra’s work was partially supported by NSF grant DMS-2210662. Kshitij Khare’s work was partially supported by NSF grant DMS-2410677.
[…]